16. Consider the titration of a 7.5 mL sample of 0.106 M HC2H3O2with 0.146 M NaOH. Ka(HC2H3O2) = 1.8x10-5

Determine the pH at one-half of the equivalence point.
Answer #1

16)

ka = 1.8 x 10^-5

pKa = - log (1.8 x 10^-5)

pKa = 4.74

millimoles of CH3COOH = 7.5 x 0.106 = 0.795

millimoles of base = 0.3975

At half - equivalence point :

CH3COOH   +   NaOH    ------------->    CH3COONa +   H2O

0.795               0.3975 0 0

0.3975 0 0.3975

pH = pKa + log [conjugate base / acid]

pH = 4.74 + log [0.3975 / 0.3975]

pH = 4.74